# RACCooN: A Versatile Instructional Video Editing Framework with Auto-Generated Narratives

### v3

@@ 3.2 P2V: User-Interactive Video Content Editing with Automatically Generated Descriptions
With the well-structured, detailed, and object-centric video description generated from the Video-to-Paragraph stage, users can ‘read’ the video details and interactively modify the content by altering the model-generated description. This approach shifts users’ focus from labor-intensive video observation to content editing. We categorize general video content editing into three important subtasks: (1) Video Object Adding: add extra objects to a video. (2) Video Object Removing: delete target objects and re-generate the object region as the background. (3) Video Object Changing: change objects’ attributes (e.g., color, textural, material). Many previous works have made great progress in video editing (Wu et al., 2024, Geyer et al., 2023, Qi et al., 2023, Zhang et al., 2023b, Fan et al., 2024) but usually focus on one of these subtasks. In this paper, we propose a unified generative model for video content editing that integrates all those crucial subtasks. Specifically, we formulate these subtasks as text-based video painting tasks and leverage a single video diffusion model for adding, removing, and changing video objects in the form of inpainting.
@@
As shown in Fig. 2 bottom, our video diffusion model processes input video $\bm{x}\in\mathbb{R}^{F\times C\times H\times W}$ with a predicted binary mask $\bm{m}^{\prime}\in\mathbb{R}^{F\times 1\times H\times W}$ targeting specific regions for modification. Following image inpainting techniques (Xie et al., 2023, Rombach et al., 2022), we apply the mask [^1] We use image grounding (Liu et al., 2023d) and video tracking models (Cheng et al., 2023) as the off-the-shelf mask predictor in inference. to the video to designate the editing region. The masked video is then encoded using a Variational Autoencoder (VAE (Kingma & Welling, 2013)) to serve as the generation condition. The model can then be informed on which video region should be edited for localized editing. Driven by the detailed description, the diffusion model can conduct diverse video editing that reflects the text prompts.
@@ 3.3 VPLM Dataset Collection and RACCooN Pipeline Training
Dataset Collection. We utilize video datasets (Majumdar et al., 2020, Gavrilyuk et al., 2018) from previous video inpainting work (Wu et al., 2024). Each raw video is accompanied by multiple inpainted versions with specific objects removed and includes binary masks of these objects. Although well-annotated with object masks and inpainted backgrounds, these datasets lack detailed descriptions of holistic video and specific local objects, hindering RACCooN’s training for producing well-structured captions for video editing. To address this, we use GPT-4V (Achiam et al., 2023) to annotate detailed video descriptions. We first re-arrange uniformly sampled video frames into a grid-image (Fan et al., 2021) and add visual prompts by numbering each frame. We then ask GPT-4V to generate detailed captions for both the entire video and key objects, in a well-structured format. Next, we train V2P and P2V stages in our framework separately (Fig. 2). In the end, RACCooN can automatically generate detailed, well-structured descriptions for raw videos and adapt these descriptions based on user updates for various video content editing tasks.
MLLM Instructional Fine-tuning. To enable the MLLM to output detailed video descriptions for content editing, we construct an instructional fine-tuning dataset based on VPLM with two video-instruction (Liu et al., 2023b) designs: (1) For object editing and removal, the MLLM generates structured video captions identifying key objects in the original video $\bm{x}$, using annotated descriptions as the learning objective. This allows users to edit videos directly from these descriptions without exhaustive analysis. (2) For object insertion, the MLLM provides not only detailed descriptions but also frame-wise placement suggestions for new objects, enhancing its utility in video editing by avoiding manual trajectory outlining. For training, we convert video object segmentation masks into bounding boxes by selecting maximal and minimal coordinates and follow the box planning strategy using LLMs (Lin et al., 2023b). We input box coordinates as a sequence of numbers and train RACCooN framework to predict these layouts given inpainted videos $\widehat{\bm{x}}$. We perform parameter-efficient fine-tuning with LoRA (Hu et al., 2022) [^2] We employ LoRA for query and value for each self-attention. on these mixed datasets with CE loss. We freeze the visual encoder and LLM backbone, updating the projector, LoRA, and LLM head.
@@ 4 Experimental Results
4.1 Experiment Setup
Tasks & Datasets: We evaluate our RACCooN framework on diverse video datasets across tasks, including video captioning (YouCook2 (Zhou et al., 2018a), VPLM), text-based video content editing (DAVIS (Pont-Tuset et al., 2017), VPLM), and conditional video generation (ActivityNet (Krishna et al., 2017), YouCook2 (Zhou et al., 2018a), UCF101 (Soomro et al., 2012)).
Metrics: For each task, we evaluate our approach with various metrics. (1) Video Caption: following previous works (Yang et al., 2023, Zhou et al., 2018b), we conduct a comprehensive human evaluation and adopt general metrics for our long video descriptions, including SPICE (Anderson et al., 2016), BLEU-4 (Vedantam et al., 2015), and CIDEr (Vedantam et al., 2015). (2) Video Object Layout Planning: following the prior work (Lin et al., 2023b), we evaluate the framework for object layout planning by bounding box IoU, FVD (Unterthiner et al., 2019), and CLIP-score (Radford et al., 2021). (3) Text-based Video Content Editing: following prior works (Geyer et al., 2023, Ceylan et al., 2023, Yang et al., 2024), we evaluate the framework by CLIP-Text, CLIP-Frame, Qedit (Yang et al., 2024), and SSIM (Hore & Ziou, 2010). (4) Conditional Video Generation: we measure FVD (Unterthiner et al., 2019), CLIP-Score (Radford et al., 2021), and SSIM (Hore & Ziou, 2010).
Implementation Details: In V2P generation, we set $k=[20,25]$ and $v=[5,6]$ for superpixel clustering. We use CLIP-L/14@336 (Radford et al., 2021) as the image encoder and Vicuna-1.5 (Zheng et al., 2024) as the LLM. Our P2V model is started from StableDiffusion-2.0-Inpainting (Rombach et al., 2022). We split the VPLM datasets into train and test sets, with the test set containing 50 unique video-paragraph pairs (for V2P) and 180 mask-object-description triples (for P2V). We manually annotate the editing prompts for the object-changing subtask. We quantitatively compare RACCooN and other baselines on the VLPM test set. To focus on generation results rather than grounding ability, we apply the same ground truth masks and captions to all methods for P2V evaluation. See the Appendix for more details on datasets, metrics, implementations, ablations, and qualitative analysis.
Table 1: Results of Single Object Prediction on VPLM test set. Metrics are abbreviated:S: SPICE, B: BLEU-4, C: CIDEr.
@@ 4.1 Experiment Setup
Figure 4: Qualitative V2P example of our RACCooN on Sora video.
4.2 Video-to-Paragraph Generation
Video-Paragraph Alignment. We conducted a quantitative evaluation of our proposed RACCooN framework’s video-to-paragraph generation capabilities, comparing it against strong baselines with a focus on object-centric captioning and object layout planning. The results, summarized in Tab. 2, show that open-source video-LLMs (e.g., PG-VL, Video-Chat) which have smaller LLMs ($<13B$ parameters), struggle with object-centric captioning and usually fail to generate layout planning. This is primarily due to their lack of instructional fine-tuning and insufficient video detail modeling without multi-granular pooling. In contrast, our RACCooN framework demonstrates superior performance in both object-centric captioning and complex object layout planning, benefiting from the instructional tuning on our VPLM dataset. Additionally, our method achieves competitive performance with proprietary MLLMs (e.g., Gemini 1.5 Pro, GPT-4o) in key object captioning and layout planning, demonstrating its superior instruction following and generation quality.
Figure 5: Qualitative Comparison between RACCooN and other baselines. Baseline names are abbreviated: VC: VideoComposer, I-A: Inpainting Anything, TF: TokenFlow. We underlined visual details in our caption. More visualizations are in Appendix.
@@ 4.2 Video-to-Paragraph Generation
LOVECon (Liao & Deng, 2023) 29.36 94.77 1.29 1319.51 60.40 17.78 1433.12 58.51 17.35
@@
Human Evaluation & Qualitative Examples. We conducted a human evaluation to compare our auto-generated captions with those from a strong baseline and human annotations on ten randomly selected YouCook2 videos (each three to five minutes long with multiple scenes and complex viewpoints). Five evaluators rated these based on Logic Fluency, Language Fluency, Video Summary, and Video Details (details in Appendix). The average scores for each criterion and their overall mean are illustrated in Tab. 2. Our method significantly surpassed both the PG-VL-generated and ground truth captions in all metrics, showing a 4.9%p and 21.8%p absolute improvement respectively, and matched the ground truth in capturing Video Details with a 9.4%p enhancement over the baseline, highlighting RACCooN’s superior capability in capturing video details. We additionally visualize descriptions generated by our RACCooN. We use a well-known generated video from the Sora (openai, 2024) generated demo example. As shown in Fig. 4, it demonstrates our model’s robust capability to auto-describe complex video content without human textual input.
@@ 4.3 Instructional Video Editing with RACCooN
Quantitative Evaluation. As shown in Tab. 3, we quantitatively compare the video editing ability of RACCooN with strong video editing models based on inpainting or DDIM-inversion (Hertz et al., 2022) across three object-centric video content editing subtasks: object changing, removal, and adding. In general, RACCooN outperforms all baselines across 9 metrics. For object changing, RACCooN outperforms the best-performing baseline by 0.8% on CLIP-T, indicating better video-text alignment while maintaining temporal consistency, as demonstrated by comparable CLIP-F and Qedit scores. Note that LGVI is not designed to alter video attributes and tends to preserve video content with marginal change (i.e., identical input and output videos), resulting in improved CLIP-F scores. In the object removal task, RACCooN shows significant improvements over strong baselines (relatively $+57.8\%$ FVD, $+2.5\%$ SSIM, $+9.6\%$ PSNR). Such improvements are maintained in the addition task (relatively $+41.6\%$ FVD, $+4.3\%$ PSNR). Meanwhile, some DDIM inversion-based models (e.g., TokenFlow (Geyer et al., 2023)) work well for specific tasks (change objects), but do not handle other types of editing. In contrast, our method is an all-rounder player.
We further emphasize that a simple combination of existing models cannot achieve an effective framework, leading to inferior instruction-following and editing abilities. This is evident in the degraded performance of open-source Video-LLM baselines (Tabs. 2 and 2) and other video editing models (Tab. 3). To address these limitations, we made unique novelty in technical and dataset contributions and achieved significant improvements in both video understanding and editing. This is evident in the comparison of our RACCooN with a multi-agent baseline combining a powerful open-source video reasoning framework and video diffusion models (PG-Video-LLaMA + StableDiffusion 2.0-inpainting) in Tab. 3. RACCooN outperforms the PGVL + SD 2.0-inpainting by a significant margin across all metrics and editing tasks, highlighting the effectiveness of our proposed framework.
@@
Ablation Studies. As shown in Tab. 4, we further validate the effectiveness of components by replacing detailed descriptions with short captions, and oracle masks/planning boxes with model-generated ones. In adding objects, our detailed object descriptions can benefit generation by providing accurate details, leading to improved quantitative results (relatively $+14.4\%$ FVD). We further replace GT boxes with boxes predicted by RACCooN, and still show superior performance over other baseline methods with oracle boxes in Tab. 3. It demonstrates that our V2P stage can thus automatically generate planning from a given video to eliminate users’ labor. Next, in object removal and changing, we replace the oracle masks with grounding (Liu et al., 2023d) and tracking (Cheng et al., 2023) tools generated mask, it shows marginally decrement for changing objects, and our framework still shows strong results over other baselines in Tab. 3 with oracle masks. It suggests that our RACCooN framework is effective and robust to handle diverse editing skills in a non-orcale setting (Please see the Appendix).
4.4 Enhancing Inversion-based Video Editing & Generation with RACCooN
We further validate that detailed paragraphs can benefit video generative tasks (e.g., inversion-based video editing and video generation). Our RACCooN captions can be integrated with off-shelf SoTA video editing and generation models to enhance them. See Appendix for more details.
Inversion-based Video Editing. Our RACCooN framework can significantly enhance video editing tasks. We integrated it into two SoTA methods, TokenFlow (Geyer et al., 2023) and FateZero (Qi et al., 2023), and compared their performance with different text inputs. The baseline used human-written short captions, while baseline+RACCooN used generated detailed captions. As shown in Tab. 5, integrating RACCooN significantly improved performance. FateZero+RACCooN achieved an 11.0% and 1.4% relative increase in CLIP-Text and SSIM scores, respectively. TokenFlow+RACCooN saw increases of 4.9% and 1.9%. These results indicate that detailed captions enhance text-to-video alignment and localized editing, while preserving temporal consistency and quality.
Conditional Video Generation. We also integrated RACCooN with various SoTA video generation methods, such as VideoCrafter (Chen et al., 2023) and DynamiCrafter (Xing et al., 2023). VideoCrafter was tested in a Text-to-Video setting, while DynamiCrafter was in an Image-Text-to-Video setting. We evaluated our framework on ActivityNet, YouCook2, and UCF101 datasets. As listed in Tab. 6, VideoCrafter’s generated videos often failed to align with the GT captions, especially in complex videos. In contrast, our framework improved video quality and alignment with auto-generated detailed paragraphs, achieving substantial improvements of 36.9% in FVD and 15.3% in SSIM. We noted consistent improvements in FVD, CLIP, and SSIM scores with RACCooN. It underscores our effectiveness in augmenting video generation models with automated descriptions.
Table 5: Results of Inversion-based Video Editing on DAVIS Video. Our generated paragraph can be integrated with different SoTA inversion-based video editing models (e.g. TokenFlow or Fate-Zero). attr. and ins. indicate attribute- and instance-level editing.
Model CLIP-Text CLIP-Frame SSIM
attr. ins. all attr. ins. all attr. ins. all
FateZero (Qi et al., 2023) 28.9 27.2 28.1 95.6 95.1 95.3 71.5 70.8 71.2
FateZero (Qi et al., 2023) + RACCooN 31.7 30.7 31.2 95.5 95.1 95.3 72.1 72.3 72.2
TokenFlow (Geyer et al., 2023) 31.4 29.8 30.6 94.6 94.1 94.3 57.0 56.0 56.5
TokenFlow (Geyer et al., 2023) + RACCooN 32.6 31.6 32.1 94.7 94.3 94.5 58.0 57.3 57.6
Table 6: Results of Conditional Video Generation on three datasets. RACCooN framework can be integrated with different video generation models (e.g. VideoCrafter or DynamiCrafter.
Model ActivityNet YouCook2 UCF101
[…]
VideoCrafter (Chen et al., 2023) 3743.62 9.58 22.26 4731.22 9.89 20.58 3556.06 9.81 18.32
VideoCrafter (Chen et al., 2023) + RACCooN 2357.41 10.53 24.02 3046.82 9.47 23.89 2208.90 9.60 22.32
DynamiCrafter (Xing et al., 2023) 1632.30 10.65 32.46 2059.93 11.95 37.22 1588.57 11.98 38.81
DynamiCrafter (Xing et al., 2023) + RACCooN 1536.63 10.69 32.86 1904.08 10.03 38.78 1573.27 9.76 39.83

### v4

@@ 3.2 P2V: User-Interactive Video Editing with Auto-Generated Descriptions
With the well-structured, detailed, and object-centric video description generated from the Video-to-Paragraph stage, users can ‘read’ the video details and interactively modify the content by altering the model-generated description. This approach shifts users’ focus from labor-intensive video observation to content editing. We categorize general video content editing into three important subtasks: (1) Video Object Adding: add extra objects to a video. (2) Video Object Removing: delete target objects and re-generate the object region as the background. (3) Video Object Changing: change objects’ attributes (e.g., color, textural, material). Many previous works have made great progress in video editing (Wu et al., 2024; Geyer et al., 2023; Qi et al., 2023; Zhang et al., 2023b; Fan et al., 2024), but usually focus on one of these subtasks. In this paper, we propose a unified generative model for video content editing that integrates all those crucial subtasks. Specifically, we formulate these subtasks as text-based video painting tasks and leverage a single video diffusion model for adding, removing, and changing video objects in the form of inpainting.
@@
As shown in Fig.˜2 bottom, our video diffusion model processes input video $\bm{x}\in\mathbb{R}^{F\times C\times H\times W}$ with a predicted binary mask $\bm{m}^{\prime}\in\mathbb{R}^{F\times 1\times H\times W}$ targeting specific regions for modification. Following image inpainting techniques (Xie et al., 2023; Rombach et al., 2022), we apply the mask [^1] We use image grounding (Liu et al., 2023d) and video tracking models (Cheng et al., 2023) as the off-the-shelf mask predictor in inference. to the video to designate the editing region. The masked video is then encoded using a Variational Autoencoder (VAE (Kingma and Welling, 2013)) to serve as the generation condition. The model can then be informed on which video region should be edited for localized editing. Driven by the detailed description, the diffusion model can conduct diverse video editing that reflects the text prompts.
In addition, we provide details regarding the process of adding objects in video editing. Indeed, adding objects can be considered a unique video editing task, distinct from removing objects or changing attributes. Unlike the latter scenarios, where the target objects are already present in the initial video, adding objects involves introducing entirely new elements, which necessitates a modified editing process. As illustrated in Fig.˜2, the MLLM in V2P provides not only detailed descriptions but also frame-wise placement suggestions for new objects in the form of bounding box sequences. The object insertion process in RACCooN in inference is conducted through the following steps:
1. User Edit: The user provides an instruction to add a specific object.
2. MLLM Output: The finetuned MLLM in V2P generates fine-grained video descriptions along with frame-wise bounding box suggestions for new objects. For example, “Layouts of <Obj> to be added: {Frame 1: [0.2, 0.0, 0.5, 0.7], Frame 2: [0.2, 0.1, 0.4, 0.65], …}” specifies the layout for each frame, where [x1, y1, x2, y2] represents the top-left and bottom-right corners of the bounding box, with coordinates normalized to the range $[0,1]$ (yellow box in Fig.˜2, top right).
3. Video Editing: Generate videos based on the MLLM-generated output, including the frame-wise layout of the object to be added.
@@ 3.3 VPLM Dataset Collection and RACCooN Pipeline Training
Dataset Collection. We utilize video datasets (Majumdar et al., 2020; Gavrilyuk et al., 2018) from previous video inpainting work (Wu et al., 2024). Each raw video is accompanied by multiple inpainted versions with specific objects removed and includes binary masks of these objects. Although well-annotated with object masks and inpainted backgrounds, these datasets lack detailed descriptions of holistic video and specific local objects, hindering RACCooN’s training for producing well-structured captions for video editing. To address this, we use GPT-4V (Achiam et al., 2023) to annotate detailed video descriptions. We first re-arrange uniformly sampled video frames into a grid-image (Fan et al., 2021) and add visual prompts by numbering each frame. We then ask GPT-4V to generate detailed captions for both the entire video and key objects, in a well-structured format. Next, we train V2P and P2V stages in our pipeline separately (Fig.˜2). In the end, RACCooN can automatically generate detailed, well-structured descriptions for raw videos and adapt these descriptions based on user updates for various video content editing tasks.
MLLM Instructional Fine-tuning. To enable the MLLM to output detailed video descriptions for content editing, we construct an instructional fine-tuning dataset based on VPLM with two video-instruction (Liu et al., 2023b) designs: (1) For object editing and removal, the MLLM generates structured video captions identifying key objects in the original video $\bm{x}$, using annotated descriptions as the learning objective. This allows users to edit videos directly from these descriptions without exhaustive analysis. (2) For object insertion, the MLLM provides not only detailed descriptions but also frame-wise placement suggestions for new objects, enhancing its utility in video editing by avoiding manual trajectory outlining. For training, we convert video object segmentation masks into bounding boxes by selecting maximal and minimal coordinates and follow the box planning strategy using LLMs (Lin et al., 2023b). We input box coordinates as a sequence of numbers and train RACCooN to predict these layouts given inpainted videos $\widehat{\bm{x}}$. We perform parameter-efficient fine-tuning with LoRA (Hu et al., 2022) [^2] We employ LoRA for query and value for each self-attention. on these mixed datasets with CE loss. We freeze the visual encoder and LLM backbone, updating the projector, LoRA, and LLM head.
@@ 4 Experimental Results
Table 1: Single Object Prediction on VPLM test set. Metrics indicate: S: SPICE, B: BLEU-4, C: CIDEr.
@@
LOVECon (Liao and Deng, 2023) 29.36 94.77 1.29 1319.51 60.40 17.78 1433.12 58.51 17.35
@@
Tasks & Datasets: We evaluate RACCooN on diverse video datasets across tasks, including video captioning (YouCook2 (Zhou et al., 2018a), VPLM), text-based video content editing (DAVIS (Pont-Tuset et al., 2017), VPLM), and conditional video generation (ActivityNet (Krishna et al., 2017), YouCook2 (Zhou et al., 2018a), UCF101 (Soomro et al., 2012)).
Metrics: For each task, we evaluate our approach with various metrics. (1) Video Caption: following previous works (Yang et al., 2023; Zhou et al., 2018b), we conduct a comprehensive human evaluation and adopt general metrics for our long video descriptions, including SPICE (Anderson et al., 2016), BLEU-4 (Vedantam et al., 2015), and CIDEr (Vedantam et al., 2015). (2) Video Object Layout Planning: following the prior work (Lin et al., 2023b), we evaluate RACCooN for object layout planning by bounding box IoU, FVD (Unterthiner et al., 2019), and CLIP-score (Radford et al., 2021). (3) Text-based Video Content Editing: following prior works (Geyer et al., 2023; Ceylan et al., 2023; Yang et al., 2024), we evaluate RACCooN’s video editing capabilities by CLIP-Text, CLIP-Frame, Qedit (Yang et al., 2024), and SSIM (Hore and Ziou, 2010). (4) Conditional Video Generation: we measure FVD (Unterthiner et al., 2019), CLIP-Score (Radford et al., 2021), and SSIM (Hore and Ziou, 2010). Implementation details are provided in the Appendix.
4.1 Video-to-Paragraph Generation
Video-Paragraph Alignment. We conducted a quantitative evaluation of our proposed RACCooN’s video-to-paragraph generation capabilities, comparing it against strong baselines with a focus on object-centric captioning and object layout planning. The results, summarized in Tab.˜2, show that open-source video-LLMs (e.g., PG-VL, Video-Chat), which have smaller LLMs (<13B<13B parameters), struggle with object-centric captioning and usually fail to generate layout planning. This is primarily due to their lack of instructional fine-tuning and insufficient video detail modeling without multi-granular pooling. In contrast, RACCooN demonstrates superior performance in both object-centric captioning and complex object layout planning, benefiting from the instructional tuning on our VPLM dataset. Additionally, our method achieves competitive performance with proprietary MLLMs (e.g., Gemini 1.5 Pro, GPT-4o) in key object captioning and layout planning, demonstrating its superior instruction following and generation quality.
Figure 4: Qualitative Comparison between RACCooN and other baselines. Baseline names are abbreviated: VC: VideoComposer, I-A: Inpainting Anything, TF: TokenFlow. We underlined visual details in our caption. More visualizations are in the supplementary material.
Human Evaluation & Qualitative Examples. We conducted a human evaluation to compare our auto-generated captions with those from a strong baseline and human annotations on ten randomly selected YouCook2 videos (each three to five minutes long with multiple scenes and complex viewpoints). Five evaluators rated these based on Logic Fluency, Language Fluency, Video Summary, and Video Details (details in the supplementary material). The average scores for each criterion and their overall mean are illustrated in Tab.˜2. Our method significantly surpassed both the PG-VL-generated and ground truth captions in all metrics, showing a 4.9%p and 21.8%p absolute improvement respectively, and matched the ground truth in capturing Video Details with a 9.4%p enhancement over the baseline, highlighting RACCooN’s superior capability in capturing video details. We additionally visualize descriptions generated by RACCooN using a well-known generated video from the Sora (openai, 2024) generated demo example. As shown in Fig. 4 (in the supplementary material), it demonstrates our robust capability to auto-describe complex video content without human textual input.
@@ 4.2 Instructional Video Editing with RACCooN
Quantitative Evaluation. As shown in Tab.˜3, we quantitatively compare the video editing ability of RACCooN with strong video editing models based on inpainting or DDIM-inversion (Hertz et al., 2022) across three object-centric video content editing subtasks: object changing, removal, and adding. In general, RACCooN outperforms all baselines across 9 metrics. For object changing, RACCooN outperforms the best-performing baseline by 0.8% on CLIP-T, indicating better video-text alignment while maintaining temporal consistency, as demonstrated by comparable CLIP-F and Qedit scores. Note that LGVI is not designed to alter video attributes and tends to preserve video content with marginal change (i.e., identical input and output videos), resulting in improved CLIP-F scores. In the object removal task, RACCooN shows significant improvements over strong baselines (relatively $57.8\%$ FVD, $2.5\%$ SSIM, $9.6\%$ PSNR). Such improvements are maintained in the addition task (relatively $41.6\%$ FVD, $4.3\%$ PSNR). Meanwhile, some DDIM inversion-based models (e.g., TokenFlow (Geyer et al., 2023)) work well for specific tasks (change objects), but do not handle other types of editing. In contrast, our method is an all-rounder player.
We further emphasize that a simple combination of existing models cannot achieve an effective video editing pipeline, leading to inferior instruction-following and editing abilities. This is evident in the degraded performance of open-source Video-LLM baselines (Tabs.˜2 and 2) and other video editing models (Tab.˜3). To address these limitations, we made unique novelty in technical and dataset contributions and achieved significant improvements in both video understanding and editing. This is evident in the comparison of RACCooN with a multi-agent baseline combining a powerful open-source video reasoning and video diffusion models (PG-Video-LLaMA + StableDiffusion 2.0-inpainting) in Tab.˜3. RACCooN outperforms the PGVL + SD 2.0-inpainting by a significant margin across all metrics and editing tasks, highlighting the effectiveness of our proposed editing method.
@@
Ablation Studies. As shown in Tab.˜4, we further validate the effectiveness of components by replacing detailed descriptions with short captions, and oracle masks/planning boxes with model-generated ones. In adding objects, our detailed object descriptions can benefit generation by providing accurate details, leading to improved quantitative results (relatively $14.4\%$ FVD). We further replace GT boxes with boxes predicted by RACCooN, and still show superior performance over other baseline methods with oracle boxes in Tab.˜3. It demonstrates that our V2P stage can thus automatically generate planning from a given video to eliminate users’ labor. Next, in object removal and changing, we replace the oracle masks with grounding (Liu et al., 2023d) and tracking (Cheng et al., 2023) tools generated mask, which shows a marginal decrement for changing objects, and RACCooN still shows strong results over other baselines in Tab.˜3 with oracle masks. It suggests that RACCooN is effective and robust to handle diverse editing skills in a non-orcale setting (See the appendix).